OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. -2x2 - 5x + 5 = 0

OpenStudy (anonymous):

\[-b \pm( \sqrt{b ^{2}-4ac})/2a\]

OpenStudy (anonymous):

First determine what your a, b, and c terms are and then plug it in to the formula.

OpenStudy (anonymous):

\[5\pm \sqrt{25-4(-2)5)}/2(-2)\] \[5\pm \sqrt{25+40}/-4\]

OpenStudy (anonymous):

\[5\pm \sqrt{65}/-4\]